# Quantum-vacuum-induced topological edge polaritons

### v1

@@ II.2 Intuitive insight at special parameter values
It is worth noting that a function of the form $x^{n}/\sqrt{n!}$ converges to 00 as $n\rightarrow\infty$ regardless of $x$. As such, $|0_{L}\rangle$ and $|0_{R}\rangle$ are normalizable for any finite value of $E_{Z}$ (provided $\Delta\neq 0$). This demonstrates the robustness of the system’s edge modes against the presence of the perturbation $E_{Z}$ that connects all spin-photon sectors.
Case ii: $E_{Z}=0$ and $J_{1}\neq 0$
In this case, $\ell=n+s/2$ is a good quantum number with which the Hamiltonian can be written as
where $\langle\ell|=\left(\langle-,0,X,j|\right)$ for $\ell=-1/2$ and $\langle\ell|=\left(\langle+,n,X,j|,\langle-,n+1,X,j|\right)$ for $\ell=n+1/2$, $X=A,B$, and $j=1,2,\cdots,N$,
$\mathcal{H}_{\rm eff,\ell=-1/2}$ $=$ $J_{1}+J_{2}\cos(\hat{k})\sigma_{x}+J_{2}\sin(\hat{k})\sigma_{y},$
$\mathcal{H}_{\rm eff,\ell\neq-1/2}$ $=$ $(J_{1}+J_{2}\cos(\hat{k}))\sigma_{x}+J_{2}\sin(\hat{k})\sigma_{y}$ (7)
$+\sqrt{\ell+1/2}\Delta\tau_{x}\sigma_{x}+\frac{\omega}{2}\tau_{z},$
we have suppressed unimportant constant terms, $\hat{k}$ is the quasimomentum operator such that $e^{\pm\mathrm{i}\hat{k}}$ shifts the lattice site ($j$) by one unit, $\sigma$’s and $\tau$’s are respectively the Pauli matrices acting on the sublattice ($A,B$) and $-\ell$ ($(+,n),(-,n+1)$) degrees of freedom.
$\mathcal{H}_{\rm eff,\ell=-1/2}$, which corresponds to the $(-,0)$ spin-photon sector, remains mathematically identical to the paradigmatic SSH model and thus continues to support a pair of zero modes as long as $J_{1}<J_{2}$. The remaining pairs of $(-,n+1)$ and $(+,n)$ spin-photon sectors are effectively described by $\mathcal{H}_{\rm eff,\ell\neq-1/2}$, which also satisfies the chiral symmetry $\mathcal{C}^{\dagger}\mathcal{H}_{\rm eff,\ell}(\hat{k})\mathcal{C}=-\mathcal{H}_{\rm eff,\ell}(\hat{k})$, particle-hole symmetry $\mathcal{P}^{\dagger}\mathcal{H}_{\rm eff,\ell}(\hat{k})\mathcal{P}=-\mathcal{H}_{\rm eff,\ell}(-\hat{k})$, and time-reversal symmetry $\mathcal{T}^{\dagger}\mathcal{H}_{\rm eff,\ell}(\hat{k})\mathcal{T}=\mathcal{H}_{\rm eff,\ell}(-\hat{k})$, but under different operators $\mathcal{C}=\tau_{x}\sigma_{z}$, $\mathcal{P}=\mathcal{K}\tau_{x}\sigma_{z}$, and $\mathcal{T}=\mathcal{K}$ respectively, where $\mathcal{K}$ is the complex conjugation operator. Therefore, all $\mathcal{H}_{\rm eff,\ell}$ belongs to the BDI class according to the Altland-Zirnbauer classification AZ97 , the topology of which is characterized by a $\mathbb{Z}$ invariant. Unlike $\mathcal{H}_{\rm eff,\ell=-1/2}$, however, $\mathcal{H}_{\rm eff,\ell\neq-1/2}$ does not exhibit a topological phase transition at $J_{1}=J_{2}$. Indeed, as demonstrated in Fig. 2(a), no gap closing is observed in the energy spectrum of $\mathcal{H}_{\rm eff,\ell\neq-1/2}$ at $J_{1}=J_{2}$. Consequently, it remains topologically trivial and does not support edge modes at all values of $J_{1}$.
where
which is defined under the assumption of periodic boundary conditions (PBC) such that $\hat{k}\rightarrow k\in[-\pi,\pi)$ is a good quantum number. We numerically compute the winding number $w_{+}$ at four different $\ell$ sectors and a fixed $J_{1}<J_{2}$ in Fig. 2(b). In all cases, it is clearly observed that $w_{+}$ is practically 00 for all $\Delta$ values under consideration. While not shown in the figure, The same results are also observed for $w_{-}$. Our winding number calculations thus confirm the absence of topological edge modes at $\ell\neq-1/2$. Therefore, the system only supports a pair of topological edge modes under the quantum vacuum sector of the Hamiltonian, i.e., as described by $\mathcal{H}_{\rm eff,\ell=-1/2}$.
Figure 2: (a) The energy spectrum of Eq. (7) at varying $J_{1}$ for different values of $\ell$ (energy shift of $\ell\omega$ has been reintroduced to account for the suppressed constant terms). (b) The winding number $w_{+}$ as a function of the light-matter coupling $\Delta$. In both panels, the parameter values are set at $J_{2}=0.5$, $\omega=2\pi$, (a) $\Delta=10$, and (b) $J_{1}=0.1$
@@ II.3 The emergence of topological edge modes at general parameter values
In Figs. 3(a) and (b), we show the few energy levels closest to $\langle a^{\dagger}a\rangle=0$ at varying intra-site hopping and magnetic field strength respectively. There, left- and right-edge-localized eigenstates could be clearly identified inside the gap between the two bulk energy levels closest to $E=0$. As demonstrated in Fig. 3(c), such edge states are absent at higher energy levels that correspond to larger $\langle a^{\dagger}a\rangle$. This serves as strong evidence that the obtained edge states are indeed induced by the quantum vacuum.
@@
where $\psi_{k}=(c_{-,A,k},c_{+,A,k},c_{-,B,k},c_{+,B,k})^{T}$, $c_{\pm,S,k}=\frac{1}{\sqrt{N}}\sum_{j}c_{\pm,S,j}e^{ijk}$ is the fermionic operator in the momentum space,
@@
is the corresponding momentum space Hamiltonian, $\sigma$’s and $\zeta$’s are respectively the Pauli matrices acting on the sublattice and spin-photon degrees of freedom, and $\zeta_{\pm}=\zeta_{x}\pm\mathrm{i}\zeta_{y}$. Upon diagonalizing $\mathcal{H}_{k}$ as a function of $k$, we found that the two energy bands closest to $E=0$ are gapped, whilst the other pairs of energy bands are gapless (see the left panel of Fig. 3(d)). As shown in the right panel of Fig. 3(d), other higher energy bands are also found to be gapless. This further confirms that topological edge states, if exist, could only emerge in the gap between the two bands closest to the quantum vacuum sector.
Finally, in Figs. 3(e) and (f), we plot the photon number expectation value $\langle a^{\dagger}a\rangle$ associated with the edge states observed in panels (b) and (a) respectively. In both cases, we found that $\langle a^{\dagger}a\rangle\approx 0$ for all parameter values under consideration, as expected from the fact that such edge states exist near the quantum vacuum condition. That $\langle a^{\dagger}a\rangle\approx 0$ increases with $E_{Z}$ in Fig. 3(e) is attributed to the fact that the magnetic field term couples the $(-,0)$ spin-photon sector with the other higher photon sectors. Nevertheless, as $\langle a^{\dagger}a\rangle$ remains close to zero even at moderate values of $E_{Z}$, such edge states remain localized near the $a^{\dagger}a=0$ photon sector, thus confirming their topological nature in the same spirit as the robust localization of ordinary topological edge states against considerable perturbations.
Figure 3: (a,b,c) The energy spectrum of Eq. (1) for $N=20$ at (a,c) varying $J_{1}$ and fixed $E_{Z}=1.5$, (b) varying $E_{Z}$ and fixed $J_{1}=0.1$. (d) The corresponding momentum space energy band structure of the system in a closed geometry at a fixed $J_{1}=0.1$ and $E_{Z}=1.5$, plotted around two energy references ($E=0$ and $E=3$) for clarity. (e,f) The photon number expectation of the edge states at varying (e) $E_{Z}$ and (f) $J_{1}$. The left-edge-(right-edge-)localized eigenstates are marked in green (red) respectively. In all panels, the other parameters are set as $J_{2}=0.5$, $\Delta=10$, $\omega=2\pi$, and the Fock space is truncated to contain up to $N_{p}=20$ photon sectors.
@@ III Absence of topological edge modes in the classical limit
where $\mathcal{I}_{f}$ is the infinite dimensional identity acting on the frequency space. By numerically calculating $w_{\rm Flo,\pm}$ over a large range of $\Delta$ values in Fig. 4(d), we verify that the time-periodic Hamiltonian is indeed topologically trivial, which is also consistent with the absence of edge states in Fig. 4(a,b).
@@ IV.1 Robustness against spatial disorder
One attractive feature of topological edge modes is their resilience against perturbations and system imperfections. To verify that our quantum-vacuum-induced edge polaritons also display such robustness, we consider the most common type of system imperfections, i.e., the presence of spatial disorder, which in our system could be simulated by making the values of all system parameters $J_{1}$, $J_{2}$, $E_{Z}$, and $\Delta$ site dependent. Specifically, for each lattice site $j$, we draw the value of the parameter $P\in\left\{J_{1},J_{2},E_{Z},\Delta\right\}$ from the uniform distribution $\left[\overline{P}-\delta P/\overline{P},\overline{P}+\delta P/\overline{P}\right]$.
Our results are summarized in Fig. 5(a-c). As expected, the obtained quantum-vacuum-induced edge polaritons (the in-gap flat line at $E\approx 0$) are evidently robust against moderate spatial disorders (up to $50\%$ disorder in $J_{1}$, $J_{2}$, and $E_{Z}$ and $5\%$ disorder in $\Delta$). At large enough disorder, the band gap around zero energy closes, thus causing any edge modes to either disappear or lose their topological robustness. The smaller resistance to disorder in $\Delta$ as compared with disorder in the other parameters is attributed to the fact that $\Delta$ is one order of magnitude larger than the other parameters; consequently, $5\%$ disorder in $\Delta$ leads to terms that are of the same order of magnitude as $50\%$ disorder in the other parameters.
@@
In Fig. 5(d), we further compute and present the disorder-averaged “wave function” profiles associated with the two quantum-vacuum-induced edge polaritons. There, we have defined the “wave function” squared as
$|\psi|^{2}(4j+2s+1+S)=\sum_{n=0}^{N_{P}}|\langle s,n,S,j|\psi\rangle|^{2},$ (22)
@@
where $s=\pm\frac{1}{2}$ is the spin state, $S=0$ ($S=1$) for sublattice A (B), $N_{p}$ represents the truncation in the Fock space to enable numerical computations, and $|\psi\rangle$ is the eigenstate corresponding to the quantum-vacuum-induced edge polariton under consideration. By closely inspecting the insets of Fig. 6(d), it is observed that the edge polariton localized on the left end has the largest support on the first unit cell at sublattice $A$ with spin --, whilst the edge polariton localized on the right end has the largest support on the last unit cell at sublattice $B$ with spin --. Remarkably, this is consistent with the idealized schematics shown in Fig. 1(a) even though the system is far away from the ideal case since $J_{1}$ and $E_{Z}$ are both nonzero, and all parameters becomes site-dependent due to the presence of disorder.
Finally, it is worth noting that the emergence of small but still significant peaks near the left (right) edge for the right-edge-(left-edge-)localized polariton is simply attributed to a numerical artefact arising from the disorder averaging process. Specifically, due to the degeneracy between the two types of edge polaritons, some left-edge-localized polaritons from different disorder realizations might be included in the disorder average computation of the right-edge-localized polariton and vice versa. Nevertheless, there remains exactly two edge polaritons near zero energy as is consistent with the topology of the system.
@@
Figure 5: (a,b,c) The disorder-averaged energy spectrum of Eq. (1) for $N=16$ under OBC. Spatial disorders are considered in (a) the hopping parameters $J_{1}$ and $J_{2}$ only with $\delta J_{1}=\delta J_{2}=\delta J$, (b) the magnetic field amplitude $E_{Z}$ only, and (c) photon-fermion interaction strength $\Delta$ only. (d) The disorder-averaged “wave function” profiles (see main text for definition) associated with quantum-vacuum-induced edge polaritons in the presence disorder on all parameters with $\delta J_{1}=\delta J_{2}=\delta E_{Z}=10\delta\Delta=0.1$ (the insets depict the zoomed-in version of each edge polariton that highlights the location of its largest peak). In all panels, $50$ disorder realizations are considered, $\omega=2\pi$ is taken, the Fock space is truncated to contain up to $N_{p}=16$ photon sectors, and the mean value of each parameter is taken as $\overline{J}_{1}=0.1$, $\overline{J}_{2}=0.5$, $\overline{\Delta}=10$, and $\overline{E}_{Z}=1.5$.
@@ IV.2 Robustness against counterrotating term effect
The TC model typically describes some physical systems under the assumption of rotating wave approximation (RWA). It amounts to approximating light-matter interaction of the form $(a+a^{\dagger})s_{x}$, which couples a particle’s spin with the electric field, by a theoretically simpler form $as_{+}+a^{\dagger}s_{-}$, where $\sigma_{\pm}=s_{x}+\mathrm{i}s_{y}$ Tavis68 . While RWA often serves as a good approximation to some physical systems, there are also instances in which RWA yields “incorrect” phenomena which are otherwise absent when the exact model is considered Wang73 ; Rzazewski75 ; Larson12 .
@@
In view of the above, we investigate the effect of the “counterrotating” interaction term on the formation of quantum-vacuum-induced edge polaritons in our model. To this end, we add
$H_{\rm int}^{\prime}=\sum_{j=1}^{N}\delta\left\{a(c^{\dagger}_{-,B,j}c_{+,A,j}+c^{\dagger}_{-,A,j}c_{+,B,j})\right\}$ (23)
to Eq. (1). There, we have introduced a new parameter $\delta$ to describe the strength of the counterrotating term. As demonstrated in Fig. 6(a), the above obtained quantum-vacuum-induced edge polaritons persist over a significant range of $\delta$ values. At $\delta/\Delta\sim 1/4$, the two center bands corresponding to the zero photon sector closes, thus signifying a topological phase transition to a topologically trivial regime that does not support quantum-vacuum-induced edge polaritons. In Fig. 6(b), we find that the mean photon number associated with the quantum-vacuum-induced edge polaritons found in Fig. 6(a) is very close to zero at small to moderate values of $\delta$, consistent with the fact that they are close to being in a vacuum photon state. Near the topological phase transition, the mean photon number of the quantum-vacuum-induced edge polaritons increases exponentially with $\delta$ before they disappear completely.
Under PBC, the momentum space energy bands of the system could be computed to uncover additional properties of the system in the presence of $\delta$. In particular, Fig. 6(c) reveals a significant energy gap between the two middle bands, which are associated with the lowest (almost zero) mean photon number. Such an energy gap enables the topological protection of the resulting edge states, i.e., the quantum-vacuum-induced edge polaritons, in the corresponding open chain setting. By contrast, the other pairs of energies shown in Fig. 6(c), which correspond to larger mean photon numbers, are gapless. This in turn prevents the existence of additional topologically protected edge states in the system when OBC are applied.
For completeness, in Fig. 6(d), we present the “wave function” profiles associated with the two quantum-vacuum-induced edge polaritons using Eq. (22). Note that both profiles look qualitatively similar to those of Fig. 5(d) despite the incorporation of a moderate counterrotating interaction parameter $\delta$ instead of spatial disorder. This result further highlights the topological nature of our quantum-vacuum-induced edge polaritons.
@@
Figure 6: (a) The energy spectrum of Eq. (1) for $N=20$ under OBC at varying counterrotating interaction amplitude $\delta$. The overlapping green and red marks highlight the quantum-vacuum-induced edge polaritons. (b) The mean photon number for the quantum-vacuum-induced edge polaritons. (c) The corresponding momentum space energy bands near $E=0$ at a fixed $\delta=1$. (d) The “wave function” profiles associated with quantum-vacuum-induced edge polaritons at a fixed $\delta=1$ (the insets depict the zoomed-in version of each edge polariton that highlights the location of its largest peak). In all panels, the system parameters are taken as $J_{1}=0.1$, $J_{2}=0.5$, $\Delta=10$, $E_{Z}=1.5$, $\omega=2\pi$, and the Fock space is truncated to contain up to $N_{p}=20$ photon sectors.

### v2

@@ II.2 Intuitive insight at special parameter values
It is worth noting that a function of the form $x^{n}/\sqrt{n!}$ converges to 00 as $n\rightarrow\infty$ regardless of $x$. As such, $|0_{L}\rangle$ and $|0_{R}\rangle$ are normalizable for any finite value of $E_{Z}$ (provided $\Delta\neq 0$). This demonstrates the robustness of the system’s edge modes against the presence of the perturbation $E_{Z}$ that connects all atom-photon sectors.
@@ II.3 The emergence of topological edge modes at general parameter values
In Figs. 3(a) and (b), we show the few energy levels closest to $\langle a^{\dagger}a\rangle=0$ at varying intra-site hopping and magnetic field strength respectively. There, all system parameters are nonzero, which means that all atom-photon sectors are nontrivially coupled. Nevertheless, left- and right-edge-localized eigenstates could be clearly identified inside the gap between the two bulk energy levels closest to $E=0$. As demonstrated in Fig. 3(c), such edge states are absent at higher energy levels that correspond to larger $\langle a^{\dagger}a\rangle$. This serves as strong evidence that the obtained edge states are indeed protected by the quantum vacuum. Moroever, it should be emphasized that while such edge states are generally no longer found at zero energy, they remain degenerate and are adiabatically connected to the zero energy edge states in the special case of $E_{Z}=0$, which are therefore topological in nature.
@@
where $\psi_{k}=(c_{-,A,k},c_{+,A,k},c_{-,B,k},c_{+,B,k})^{T}$, $c_{\pm,S,k}=\frac{1}{\sqrt{N}}\sum_{j}c_{\pm,S,j}e^{ijk}$ is the particle annihilation operator in the momentum space,
@@
is the corresponding momentum space Hamiltonian, $\sigma$’s and $\zeta$’s are respectively the Pauli matrices acting on the sublattice and internal degrees of freedom, and $\zeta_{\pm}=\zeta_{x}\pm\mathrm{i}\zeta_{y}$. Upon diagonalizing $\mathcal{H}_{k}$ as a function of $k$, we found that the two energy bands closest to $E=0$ are gapped (see the left panel of Fig. 3(d)), whilst the other pairs of energy bands are gapless (see both panels of Fig. 3(d) for some representative gapless bands). This further confirms that topological edge states, if exist, could only emerge in the gap between the two bands closest to the quantum vacuum sector.
Finally, in Figs. 3(e) and (f), we plot the photon number expectation value $\langle a^{\dagger}a\rangle$ associated with the edge states observed in panels (b) and (a) respectively. In both cases, we found that $\langle a^{\dagger}a\rangle\approx 0$ for all parameter values under consideration, as expected from the fact that such edge states exist near the quantum vacuum sector. That $\langle a^{\dagger}a\rangle\approx 0$ increases with $E_{Z}$ in Fig. 3(e) is attributed to the fact that the magnetic field term couples the $(-,0)$ atom-photon sector with the other higher photon sectors. Nevertheless, as $\langle a^{\dagger}a\rangle$ remains close to zero even at moderate values of $E_{Z}$, such edge states remain localized near the $a^{\dagger}a=0$ photon sector, thus confirming their topological nature in the same spirit as the robust localization of ordinary topological edge states near a physical edge against considerable perturbations. Due to the crucial role of quantum vacuum as an artificial edge that facilitates the above obtained edge states, the latter shall be referred to as quantum-vacuum-protected edge polaritons in the remainder of this paper.
Figure 3: (a,b,c) The energy spectrum of Eq. (1) for $N=20$ at (a,c) varying $J_{1}$ and fixed $E_{Z}=1.5$, (b) varying $E_{Z}$ and fixed $J_{1}=0.1$. (d) The corresponding momentum space energy band structure of the system in a closed geometry at a fixed $J_{1}=0.1$ and $E_{Z}=1.5$, plotted around two energy references ($E=0$ and $E=3$) for clarity. (e,f) The photon number expectation values of the edge states at varying (e) $E_{Z}$ and (f) $J_{1}$. The left-edge-(right-edge-)localized eigenstates are marked in green (red) respectively. In all panels, the other parameters are set as $J_{2}=0.5$, $\Delta=10$, $\omega=2\pi$, and the Fock space is truncated to contain up to $N_{p}=20$ photon sectors.
@@ III Absence of topological edge modes in the classical limit
where $\mathcal{I}_{f}$ is the infinite dimensional identity matrix acting on the frequency space. By numerically calculating $w_{\rm Flo,\pm}$ over a large range of $\Delta$ values in Fig. 4(d), we verify that the time-periodic Hamiltonian is indeed topologically trivial, which is also consistent with the absence of edge states in Fig. 4(a,b).
@@ IV.1 Robustness against spatial disorder
One attractive feature of topological edge modes is their resilience against perturbations and system imperfections. To verify that our quantum-vacuum-protected edge polaritons also display such robustness, we consider the most common type of system imperfections, i.e., the presence of spatial disorder, which in our system could be simulated by making the values of some or all system parameters $J_{1}$, $J_{2}$, $E_{Z}$, and $\Delta$ site dependent. Specifically, for each lattice site $j$, we draw the value of the parameter $P\in\left\{J_{1},J_{2},E_{Z},\Delta\right\}$ from the uniform distribution $\left[\overline{P}-\delta P\overline{P},\overline{P}+\delta P\overline{P}\right]$.
Our results are summarized in Fig. 5. To understand the impact of disorder on each parameter individually, disorder is only applied on one parameter in panels (a)-(c), i.e., hopping parameters alone in panel (a), magnetic field strength alone in panel (b), and light-matter interaction alone in panel (c). For completeness, we also consider the effect of disorder on all parameters simultaneously in panel (d), which is particularly relevant in experiments where fine-tuning all parameters exactly to a single set of values is typically impossible. As expected, the obtained quantum-vacuum-protected edge polaritons (the in-gap flat line at almost zero energy) are evidently robust against moderate spatial disorders in all scenarios (up to $50\%$ disorder in $J_{1}$, $J_{2}$, and/or $E_{Z}$ and $5\%$ disorder in $\Delta$). At large enough disorder, the band gap around zero energy closes, thus causing any edge modes to either disappear or lose their topological robustness.
In Fig. 5(a,b), it is observed that the disorder-induced band gap closing occurs at a slightly smaller disorder strength for disorder in hopping parameters only as compared to disorder in the magnetic field. This could be attributed to the fact that the hopping parameters directly affect the topological phase of each SSH chain, whereas the magnetic field only serves as a perturbation that generally preserves the system’s topology. Finally, the smaller resistance to disorder in $\Delta$ as compared with disorder in the other parameters is attributed to the fact that $\Delta$ is one order of magnitude larger than the other parameters; consequently, $5\%$ disorder in $\Delta$ leads to terms that are of the same order of magnitude as $50\%$ disorder in the other parameters.
@@
In Fig. 6, we further compute and present the disorder-averaged “wave function” profiles associated with the two quantum-vacuum-protected edge polaritons for each disorder scenario considered in Fig. 5. There, we have defined the “wave function” squared as
$|\psi|^{2}(4j+s+S+1)=\sum_{n=0}^{N_{P}}|\langle s,n,S,j|\psi\rangle|^{2},$ (22)
@@
where $s=\pm 1$ is the internal level state, $S=0$ ($S=1$) for sublattice A (B), $N_{p}$ represents the truncation in the Fock space to enable numerical computations, and $|\psi\rangle$ is the eigenstate corresponding to the quantum-vacuum-protected edge polariton under consideration. The combination $4j+2s+1+S$ in Eq. (22) is chosen so that every $(j,s,S)$ is mapped to a unique integer, e.g., $(0,-1,0)$ is mapped to 00, $(0,-1,1)$ is mapped to $1$, $(0,1,0)$ is mapped to $2$, etc. By closely inspecting the insets of any panel in Fig. 6, it is observed that the edge polariton localized on the left end has the largest support on $04j+s+S+1=$ 0, which corresponds to the first unit cell at sublattice $A$ with internal level $-1$. Meanwhile, the edge polariton localized on the right end has the largest support on $614j+s+S+1=$ 61, which for $N=16$ corresponds to the last unit cell at sublattice $B$ with internal level $-1$. Remarkably, this is consistent with the idealized schematics shown in Fig. 1(a) even though the system is far away from the ideal case since $J_{1}$ and $E_{Z}$ are both nonzero, and all parameters become site-dependent due to the presence of disorder. Moreover, there remains exactly two edge polaritons near zero energy, which is consistent with the topology of the system.
@@
Figure 5: The disorder-averaged energy spectrum of Eq. (1) for $N=16$ under OBC. Spatial disorders are considered in (a) the hopping parameters $J_{1}$ and $J_{2}$ only with $\delta J_{1}=\delta J_{2}=\delta J$, (b) the magnetic field amplitude $E_{Z}$ only, (c) light-matter interaction strength $\Delta$ only, and (d) all parameters with $\delta J_{1}=\delta J_{2}=\delta E_{Z}=10\delta\Delta=\delta P$. In all panels, $50$ disorder realizations are considered, $\omega=2\pi$ is taken, the Fock space is truncated to contain up to $N_{p}=16$ photon sectors, and the mean value of each parameter is taken as $\overline{J}_{1}=0.1$, $\overline{J}_{2}=0.5$, $\overline{\Delta}=10$, and $\overline{E}_{Z}=1.5$.
Figure 6: The disorder-averaged “wave function” profiles (see Eq. (22) and its accompanying paragraph for the precise definition) associated with quantum-vacuum-protected edge polaritons in the presence of disorder on the parameters considered in Fig. 5. For the cases of disorder in the hopping parameters (panels (a,d)) and/or the magnetic field (panels (b,d)), the disorder strength is set at $0.1$. Meanwhile, for the cases of disorder in the interaction strength (panels (c,d)), the disorder strength is set at $0.01$. The insets depict the zoomed-in version of each edge polariton that highlights the location of its largest peak. All other parameters are the same as in Fig. 5
@@ IV.2 Robustness against counterrotating term effect
The TC model typically describes some physical systems under the assumption of rotating wave approximation (RWA). It amounts to approximating light-matter interaction of the form $(a+a^{\dagger})\sigma_{x}$, which couples a particle’s dipole moment with the electric field, by a theoretically simpler form $a\sigma_{+}+a^{\dagger}\sigma_{-}$, where $\sigma_{\pm}=\sigma_{x}\pm\mathrm{i}\sigma_{y}$ Tavis68 . While RWA often serves as a good approximation to some physical systems, there are also instances in which RWA yields “incorrect” phenomena which are otherwise absent when the exact model is considered Wang73 ; Rzazewski75 ; Larson12 .
@@
In view of the above, we investigate the effect of the “counterrotating” interaction term on the formation of quantum-vacuum-protected edge polaritons in our model. To this end, we add
$H_{\rm int}^{\prime}=\sum_{j=1}^{N}\delta\left\{a(c^{\dagger}_{-,B,j}c_{+,A,j}+c^{\dagger}_{-,A,j}c_{+,B,j})+h.c.\right\}$ (23)
to Eq. (1). There, we have introduced a new parameter $\delta$ to describe the strength of the counterrotating term. In Fig. 7(a,b), we plot the system’s energy spectrum at varying $\delta/\Delta$ while fixing $\Delta$. The quantum-vacuum-protected edge polaritons are identified as either left- or right-end localized edge modes (marked by green or red marks in Fig. 7(a,b) respectively) with almost zero mean photon number (as confirmed in Fig. 7(c)). There, we observe that quantum-vacuum-protected edge polaritons, exist over a significant range of parameter values with $\delta/\Delta\lessapprox 1/4$ or $\delta/\Delta\gtrapprox 4$. That is, quantum-vacuum-protected edge polaritons are observed not only when $\delta$ is sufficiently small, but also when $\delta$ is very large. This observation is attributed to the fact that the counterrotating interaction term also leads to the pairing between two atom-photon sectors, i.e., $(-,n)$ and $(+,n+1)$. Therefore, in the absence of the rotating term $(\Delta=0)$, the quantum vacuum sector $(+,0)$ is left uncoupled and has the potential for supporting edge modes. When $\delta$ and $\Delta$ are both nonzero, there is a competition between the different pairing mechanisms induced by both interaction terms. Consequently, quantum-vacuum-protected edge polaritons are expected to be present only when one interaction type (rotating or counterrotating) is sufficiently more dominant than the other.
That the quantum-vacuum-protected edge polaritons are present over a window of parameter values rather than at a specific value of either $\delta=0$ or $\Delta=0$ is attributed to their topological nature. The latter is further supported by the fact that the disappearance of the quantum-vacuum-protected edge polaritons is accompanied by an energy gap closing, as also seen in Fig. 7(a,b). In particular, Fig. 7(b) highlights the fact that the quantum-vacuum-protected edge polaritons indeed arise inside a small but finite energy gap between $\delta/\Delta=4$ and $\delta/\Delta=5$, which is not visible from Fig. 7(a).
In Fig. 8(a,b), we further plot the system’s energy spectrum at varying $\delta$ and a fixed $\Delta$, and vice versa. As expected, the spectra look virtually identical and the quantum-vacuum-protected edge polaritons are observed for $\delta/\Delta\lessapprox 1/4$ or $\delta/\Delta\gtrapprox 4$ (topologically nontrivial regime). In Fig. 8(c,d), we verify that the mean photon number corresponding to each quantum-vacuum-protected edge polariton is indeed very close to zero deep in the topologically nontrivial regime, consistent with the fact that such modes are close to being in a vacuum photon state. Near the topological phase transition, the mean photon numbers of the quantum-vacuum-protected edge polaritons increase exponentially before they disappear completely. This is analogous to the behavior of an ordinary topological edge state that becomes more delocalized in a physical space as it approaches a topological phase transition point.
@@
Figure 7: (a) The energy spectrum of the system for $N=20$ under OBC at varying $\delta/\Delta$ while keeping $\Delta=4$ fixed. (b) A zoomed-in view of the energy spectrum between $\delta/\Delta=4$ to $\delta/\Delta=5$. The overlapping green and red marks highlight the quantum-vacuum-protected edge polaritons localized near the system’s left and right ends respectively. (c) The mean photon number for the obtained quantum-vacuum-protected edge polaritons. In all panels, the system parameters are taken as $J_{1}=0.1$, $J_{2}=0.5$, $E_{Z}=1.5$, $\omega=2\pi$, and the Fock space is truncated to contain up to $N_{p}=20$ photon sectors.
Under PBC, the momentum space energy bands could be computed to uncover additional properties of the system in the presence of $\delta$. In particular, Fig. 8(c) reveals a significant energy gap between the two middle bands, which are associated with the lowest (almost zero) mean photon number. Such an energy gap enables the topological protection of the resulting edge states, i.e., the quantum-vacuum-protected edge polaritons, in the corresponding open chain setting. By contrast, the other pairs of energies shown in Fig. 8(c), which correspond to larger mean photon numbers, are gapless. This in turn prevents the existence of additional topologically protected edge states in the system when OBC are applied.
For completeness, in Fig. 8(d), we present the “wave function” profiles associated with the two quantum-vacuum-protected edge polaritons using Eq. (22). Note that both profiles look qualitatively similar to those of Fig. 6 despite the incorporation of a moderate counterrotating interaction parameter $\delta$ instead of spatial disorder. This result further highlights the topological nature of our quantum-vacuum-protected edge polaritons.
We close this section by highlighting the fact that since our system is in the topologically trivial regime at $\delta=\Delta$, our system cannot simply be understood as an RWA description of some physical system that only contains the usual dipole-electric-field type of light-matter interaction (corresponding to the ideal $\delta=0$). In this case, an additional type of light-matter interaction term should be added to the system to generate a difference between the rotating and counterrotating interaction amplitudes without imposing RWA. While obtaining such an additional interaction term is a nontrivial task, it is not impossible given the rapid progress in circuit quantum electrodynamics and artificial atoms technology Fink09 . Finally, it is also worth mentioning that due to the topological nature of our system, any fine tuning in generating the required interaction term is unnecessary.
Figure 8: (a,b) The energy spectrum of Eq. (1) for $N=20$ under OBC at (a) varying counterrotating interaction amplitude $\delta$ and a fixed $\Delta=10$ and (b) varying rotating interaction amplitude $\Delta$ and a fixed $\delta=10$. The overlapping green and red marks highlight the quantum-vacuum-protected edge polaritons localized near the system’s left and right ends respectively. (c,d) The mean photon number for the quantum-vacuum-protected edge polaritons associated with (a,b) respectively. (e) The corresponding momentum space energy bands near $E=0$ at a fixed $\delta=1$ and $\Delta=10$. (f) The “wave function” profiles associated with quantum-vacuum-protected edge polaritons at a fixed $\delta=1$ and $\Delta=10$ (the insets depict the zoomed-in version of each edge polariton that highlights the location of its largest peak). All other parameters are chosen to be the same as in Fig. 7.